OpenStudy (anonymous):

Someone please help!! the sum of two numbers is 73. five times the smaller number exceeds the larget number by 149. find the numbers.

OpenStudy (anonymous):

is really this is ur question plz cross check it !!

OpenStudy (anonymous):

yes this is my question

OpenStudy (anonymous):

let the two numbers be x & y => x + y = 73 => 5x + 149 = y => 5x - y = -149 now solve, for x and y !!!!!!!!!!

OpenStudy (anonymous):

but by this u will get a negative answer

OpenStudy (anonymous):

x + y = 73 5x - y = -149 add ------------- 6x =-76 x=-76/6=-38/3=-12.67

OpenStudy (anonymous):

do you have question?

OpenStudy (anonymous):

x=-12.67 then x + y = 73 -12.67+y=73 y=73+12.67=85.67